OpenStudy (anonymous):

7 over 3 and 5 over 4=

OpenStudy (anonymous):

all of it =to

OpenStudy (paki):

7/3 * 5/4....? is this rite...?

OpenStudy (anonymous):

yes

OpenStudy (paki):

simple.... 7*5 / 3*4

OpenStudy (paki):

solve it now...

OpenStudy (anonymous):

i an have tubke solving can wock ne theow it

OpenStudy (paki):

7*5 / 3*4 = 35/12...

OpenStudy (anonymous):

ok thinks

OpenStudy (paki):

pleasure

OpenStudy (anonymous):

3/7 ÷ 4/9 = ??? 3/7 1 1/27 1 27/28

OpenStudy (anonymous):

last one @cece1902